What is the average of seven numbers, preceding 7, each decreasing by 1/2?
Answer by math-vortex(648)


Hi, there--

The Problem:
What is the average of seven numbers, preceding 7, each decreasing by 1/2?

Solution:
Start with 7 and subtract 1/2. I will use the decimal form of 1/2 because decimals display 
between on this website, and they are easier to manipulate. (1/2 = 0.5)
7 - 0.5 = 6.5

The first number is 6.5. (We don't use 7 because 7 does not precede 7.)

Repeat this process until you have 7 numbers. The numbers are
6.5, 6.0, 5.5, 5.0, 4.5, 4.0, 3.5

We find the average of a set of numbers by finding their sum and dividing it by the number 
in the set.
6.5 + 6.0 + 5.5 + 5.0 + 4.5 + 4.0 + 3.5 = 35

35 divided by 7 is 5.

What is the average of seven numbers, preceding 7, each decreasing by 1/2 is 5.

Good luck!
Mrs. Figgy